OpenStudy (anonymous):

solve for all possible x:\[{2\tan x\over1-\tan^2x}=\frac1{\sqrt3}\]

OpenStudy (turingtest):

@jamie.marie all those capital letters are a bit straining on the eyes, could you please type out questions more clearly in the future?

OpenStudy (anonymous):

Sorry i forgot to turn it off!

OpenStudy (turingtest):

solve for x:\[{2\tan x\over1-\tan^2x}=\frac1{\sqrt3}\]correct?

OpenStudy (anonymous):

yes.

OpenStudy (turingtest):

I will modify your post for you

OpenStudy (anonymous):

\[ \tan(2 x) = \frac{2\tan(x)}{1 - \tan^2(x)} \]

OpenStudy (anonymous):

You need to solve \[ \tan(2 x) =\frac 1 {\sqrt 3} \] What angle has as tangent \[\frac 1 {\sqrt 3} \]